If ux² + y² +2²,v=x+y+z₁w=xy+yz+zx, then the value of the jacobian is d(u,v,w)/d(x,y,z)


1
Expert's answer
2022-02-22T10:53:51-0500

Solution:

"u=x\u00b2 + y\u00b2 +z\u00b2,v=x+y+z,w=xy+yz+zx"

Jacobian "=\begin{vmatrix}\dfrac{\partial u}{\partial x}&\dfrac{\partial u}{\partial y}&\dfrac{\partial u}{\partial z}
\\\dfrac{\partial v}{\partial x}&\dfrac{\partial v}{\partial y}&\dfrac{\partial v}{\partial z}
\\\dfrac{\partial w}{\partial x}&\dfrac{\partial w}{\partial y}&\dfrac{\partial w}{\partial z} \end{vmatrix}"

"=\begin{vmatrix}2x&2y&2z
\\1&1&1
\\y+z&x+z&x+y\end{vmatrix}
\\=-2\begin{vmatrix}1&1&1
\\x&y&z
\\y+z&x+z&x+y\end{vmatrix}
\\=-2[y(x+y)-z(x+z)-x(x+y)+z(y+z)+x(x+z)-y(y+z)]
\\=-2[xy+y^2-xz-z^2-x^2-xy+yz+z^2+x^2+xz-y^2-yz]
\\=-2(0)
\\=0"